OpenStudy (anonymous):

1 gallon = 128 ounces 1 liter = 33.8 ounces 1 quart = 32 ounces 1 pint = 16 ounces According to the table below ,which one of the following contain the most liquid container is full? A 2 1/2 liter bottle B 3/4 gallon jug C 5 1/2 pint carton D 3 1/2 quart vase

OpenStudy (beginnersmind):

Convert all of them to ounces and you can compare directly.

OpenStudy (kash_thesmartguy):

What do you think?

OpenStudy (anonymous):

yes , I know cover but how I know which now is full?

OpenStudy (beginnersmind):

I think they want you to assume that all of them are full, and ask which one contains the most liquid :)

OpenStudy (kash_thesmartguy):

\[2.5*33.8=84.5oz\]\[0.75*128=96oz\]\[5.5*16=88oz\]\[3.5*32=112oz.\]Here, we can clearly see that D has the most liquid. D is the correct answer!

OpenStudy (kash_thesmartguy):

You get it, right? @Nancy_Lam

OpenStudy (anonymous):

ok, thanks